How can you put an equation in standard form when they give you a problem and another set of points?
ex: (2,13) 5x-3y=-16[I am not sure if that is a good example(sorry)]

Answer by Cintchr(481)

This works fine. Infact .. the equation you have given IS in standard form.
+5x-3y=-16+
Ax + By = C is Standard form